# Synchronous Consensus in Partial Synchrony

### v1

@@ 3.2. Synchronous Consensus Protocol
At the end of Phase Two every process, in the largest majority of correct processes with bounded delivery between themselves, knows the initial value and the agreement proposal of every other process in that majority. This information is sufficient for every process in the majority to decide exactly the same valid agreement value at exactly the same point of time on their individual clocks.
@@ 3.4. The Trim Rule
The protocol requires every correct process that is in a majority set to decide about the agreement value. If process $P$ is in the largest majority set $C$, then $\geq(n-f)$ processes must decide the same agreement value. Yet $P$ being in a majority set $D$, $|D|\geq(n-f)$, does not know whether $D$ is the set $C$. So, every correct process $P$ must decide by following the so called Trim Rule as the rest of the correct processes in the same position. Thus, if $D$ is the set $C$, the rest of processes in $D$ will decide exactly the same agreement value.
@@
At Phase Two the size of set C might have shrunk. Yet the number of processes in $C$ is still $\geq(n-f)$ according to the system model. So all processes in $C$ at the end of Phase Two decide the same agreement value, unless two sets of equal size satisfy the criteria for being the set $C$ and the processes in their intersection cannot decide how to apply the Trim Rule. To show that the Trim Rule cannot prevent a majority from deciding the same agreement value we have either to prove that a majority decides the same value even if Phase Two completes with two largest sets of equal size for the purposes of the Trim Rule, or to show that completion of Phase Two with with two largest sets of equal size for the purposes of the Trim Rule is infeasible.
@@ 4.2. Consensus Protocol
- Case 1A: Omission to send its own protocol message to all recipients. The effect is the same as with a process crash-fail before the broadcast. It cannot spoil anything [Lemma 5].
- Case 1B: Omission to send its own protocol message to some recipients. The effect is the same as with a process crash-fail during the broadcast. It cannot spoil anything [Lemma 5].
@@
Case 2: Delay before sending its own protocol message to some or all recipients, or before retransmission of a received protocol message as part of an indirect delivery channel.
- Case 2A: Delay before sending its own protocol message to some or all recipients and before retransmission of a received protocol message as part of an indirect delivery channel to some or all recipients. The effect is the same as with a correct process with asynchronous outgoing links before the broadcast and after the broadcast. It cannot spoil anything [Lemma 4].
- Case 2B: Delay before sending its own protocol message to some or all recipients but not before retransmission of a received protocol message as part of an indirect delivery channel to some or all recipients. The effect is the same as with a correct process operating with asynchronous outgoing links that existed before the broadcast but regained synchrony after the broadcast. It cannot spoil anything [Lemma 4].
@@
Let $n_{F}$ be the number of processes that crashed before the start of current consensus round, $n_{B}$ be the number of Byzantine processes, and $n_{Cr}$ be the number of processes that may crash during the current consensus round, so that $n_{F}+n_{B}+n_{Cr}\leq f$. Trivial are the cases where the system model does not allow more crash-failed or Byzantine processes and new asynchronous links cannot affect $|C|$: i) i) $n_{F}=f$; and ii) $(n_{F}<f)\land(|C|=(f+1))$. Non-trivial configuration of a consensus system is: $((n_{F}+n_{B}+n_{Cr})<f)\land(|C|>(f+1))\land((n_{F}+n_{B}+n_{Cr}+|C|)<n)$. It allows more processes in $C$ to crash or become Byzantine, more links to become asynchronous, and more correct processes to establish two-way bounded delivery with every process in $C$ from asynchronous links regaining synchrony. The system of Lemma 4 fits into the non-trivial configuration, except in regard to possibility for crash-failed and Byzantine processes. We have to show that adding to Lemma 4 the possibility for crash-failed and Byzantine processes cannot affect its correctness. Let $P_{Cr}$ be a process that may crash during consensus round and $P_{A}$ be a process that may drop from the largest set of processes with bounded delivery due to asynchrony of links. Considering the possible dynamic disappearing of asynchrony, Lemma 4 operates with sets $C’^{\prime}\subseteq C’\subseteq C$ of correct processes with bounded delivery within $C^{\prime\prime}$, where $|C^{\prime\prime}|\geq(n-f)$ according to the system model. $C’\subseteq C$ reflects a possibility that processes $P_{Cr}$ and $P_{A}$ might have been in $C$, but were not in $C^{\prime}$ at the end of Phase One. $C’^{\prime}\subseteq C^{\prime}$ reflects a possibility that processes $P_{Cr}$ and $P_{A}$ might have been in $C^{\prime}$, but were not in $C^{\prime\prime}$ at the end of Phase Two.
When $|C|=|C^{\prime}|+2$ and $|C^{\prime}|=|C^{\prime\prime}|+2$, Lemma 4 considers that dynamically increased number of asynchronous links caused the difference However, if it is caused by $P_{A}$ and $P_{Cr}$, this cannot affect the correctness of Lemma 4. Moreover, when $|C|=|C^{\prime}|+3$ and $|C^{\prime}|=|C^{\prime\prime}|+3$, in view of Lemma 6 about the inability of a Byzantine process $P_{B}$ to cause more harm than $P_{Cr}$, when the difference is caused by $P_{A}$, $P_{Cr}$, and $P_{B}$, the correctness of Lemma 4 would be still unaffected.
@@ 5. Summary of Findings
In pursuing the objectives of this work using up to 3-hop indirect delivery for tolerance to link faults and asynchrony, we found: 1. The number of tolerated link faults / links asynchrony per system size and the actual number of faulty processes. 2. Impossibility of two equal in size largest consensus majority sets operating with bounded delivery. 3. Impossibility of two consensus majority sets of different size with bounded delivery and intersection smaller than a majority. 4. Possibility for $f$-resilience to Byzantine faults implemented with two rounds of exchange of messages. 5. Impossibility of transient nature of links asynchrony and link faults to affect consensus safety, validity, or termination. 6. The same tolerance to max allowed process crashes regardless of whether crashes happened before or during consensus rounds. 7. A way to make leaderless consensus in partial synchrony a non-existing theoretical or practical problem. 8. Possibility for consensus algorithm in partial synchrony with ensured simultaneous validity, safety, and bounded liveness.
@@ 6. Conclusion
We demonstrated: i) implementation of $f$-resilience to Byzantine faults with only two messaging rounds; ii) leaderless consensus in partially synchronous communication environment; and iii) the conceptual and practical possibility for deterministic consensus algorithm with ensured simultaneous validity, safety, and time-bounded liveness in partial synchrony. The presented possibility for time-bounded termination is a result from circumvention of a known amount of asynchronous / faulty links per number of process faults and the ensured thereby bounded delivery of a number of consensus protocol messages that is sufficient for a valid agreement. This in turn enables the use of synchronous consensus protocols operating with no need of a leader. When a synchronous protocol operates with agreement that does not depend on the content of the initial values, the protocol does not face the eventuality of non-termination and terminates within the known time-bound. The price for these possibilities is the necessity to exchange $2*n^{3}$ messages. Compared to the authenticated Byzantine algorithm of Dolev and Strong (Dolev and Strong, 1983), which operates with $(n/2+1)*n^{2}$, our algorithm requires exchange of nearly 4 times more messages. This price seems fair for the ability of a Byzantine algorithm to tolerate a known amount of communication asynchrony and faults. However, the number of exchanged messages in this case can be misleading. The algorithm of Dolev and Strong operates with $(n/2+1)$ rounds of exchange of Initial Value messages, while our operates with one round of exchange of Initial Value messages and one round of exchange of Agreement Proposal messages, which contain only hash digests and signatures, hence their size is constant and small. For instance, a system of 11 processes and 100kB initial value data: in the case with $(n/2+1)$ rounds of exchange of initial values it transfers 66.11MB, while in the case of one round with initial values and one round with agreement proposals it transfers 99.99MB. Thus, the time-bounded Byzantine fault tolerance in partial synchrony is paid with roughly 50 percent higher communication workload than the well-known Byzantine fault tolerance in complete synchrony. The presented results are anticipated to enable non-vulnerable interoperability across data systems. The concept of the International Monetary Fund about decentralised foreign exchange markets for tokenised currencies (Adrian et al., 2023), where settlement happens simultaneously with a trade and no clearing services are needed, points out its critical role for feasibility of world-shaking innovations.

### v3

@@ 3.2. Synchronous Consensus Protocol
At the end of Phase Two every process, in the largest majority of correct processes with bounded delivery between themselves, knows the initial value and the agreement proposal of every other process in that majority. This information is sufficient for every process in the majority to decide exactly the same valid agreement value at exactly the same point of time.
@@ 3.4. The Trim Rule
The protocol requires every correct process that is in a majority set to decide about the agreement value. If process $P$ is in the largest majority set $C$, then $\geq(n-f)$ processes must decide the same agreement value. Yet $P$ being in a majority set $D$, $|D|\geq(n-f)$, does not know whether $D$ is the set $C$. So, every correct process $P$ must decide by following the so called Trim Rule as the rest of the correct processes in the same position. Thus, if $D$ is the set $C$, the rest of processes in $D$ will decide the same agreement value.
@@
At Phase Two the size of set $C$ might have shrunk but the number of processes in $C$ is still $\geq(n-f)$ according to the system model. So all processes in $C$ at the end of Phase Two decide the same agreement value, unless two sets of equal size satisfy the criteria for being the set $C$ and the processes in their intersection cannot decide how to apply the Trim Rule. We will show that this cannot happen.
@@ 4.2. Consensus Protocol
- Case 1A: Omission to send own protocol message to all recipients. The effect is the same as with a process crash-fail before the broadcast. It cannot spoil anything [Lemma 5].
- Case 1B: Omission to send own protocol message to some recipients. The effect is the same as with a process crash-fail during the broadcast. It cannot spoil anything [Lemma 5].
@@
Case 2: Delay before sending own protocol message to some or all recipients, or before retransmission of a received protocol message as part of an indirect delivery channel.
- Case 2A: Delay before sending own protocol message to some or all recipients and before retransmission of a received protocol message as part of an indirect delivery channel to some or all recipients. The effect is the same as with a correct process with asynchronous outgoing links before the broadcast and after the broadcast. It cannot spoil anything [Lemma 4].
- Case 2B: Delay before sending own protocol message to some or all recipients but not before retransmission of a received protocol message as part of an indirect delivery channel to some or all recipients. The effect is the same as with a correct process operating with asynchronous outgoing links that existed before the broadcast but regained synchrony after the broadcast. It cannot spoil anything [Lemma 4].
@@
Let $n_{F}$ be the number of processes that crashed before the start of current consensus round, $n_{B}$ be the number of Byzantine processes, and $n_{Cr}$ be the number of processes that may crash during the current consensus round, so that $n_{F}+n_{B}+n_{Cr}\leq f$. Trivial are the cases where the system model does not allow more crash-failed or Byzantine processes and new asynchronous links cannot affect $|C|$: i) i) $n_{F}=f$; and ii) $(n_{F}<f)\land(|C|=(f+1))$. Non-trivial configuration of a consensus system is: $((n_{F}+n_{B}+n_{Cr})<f)\land(|C|>(f+1))\land((n_{F}+n_{B}+n_{Cr}+|C|)<n)$. It allows more processes in $C$ to crash or become Byzantine, more links to become asynchronous, and more correct processes to establish two-way bounded delivery with every process in $C$ from asynchronous links regaining synchrony. The system of Lemma 4 fits into the non-trivial configuration, except in regard to possibility for crash-failed and Byzantine processes. We have to show that adding to Lemma 4 the possibility for crash-failed and Byzantine processes cannot affect its correctness.
Let $P_{Cr}$ be a process that may crash during consensus round and $P_{A}$ be a process that may drop from the largest set of processes with bounded delivery due to asynchrony of links. Considering the possible dynamic disappearing of asynchrony, Lemma 4 operates with sets $C’^{\prime}\subseteq C’\subseteq C$ of correct processes with bounded delivery within $C^{\prime\prime}$, where $|C^{\prime\prime}|\geq(n-f)$ according to the system model. $C’\subseteq C$ reflects a possibility that processes $P_{Cr}$ and $P_{A}$ might have been in $C$, but were not in $C^{\prime}$ at the end of Phase One. $C’^{\prime}\subseteq C^{\prime}$ reflects a possibility that processes $P_{Cr}$ and $P_{A}$ might have been in $C^{\prime}$, but were not in $C^{\prime\prime}$ at the end of Phase Two. When $|C|=|C^{\prime}|+2$ and $|C^{\prime}|=|C^{\prime\prime}|+2$, Lemma 4 considers that dynamically increased number of asynchronous links caused the difference. However, if it is caused by $P_{A}$ and $P_{Cr}$, this cannot affect the correctness of Lemma 4. Moreover, when $|C|=|C^{\prime}|+3$ and $|C^{\prime}|=|C^{\prime\prime}|+3$, in view of Lemma 6 about the inability of a Byzantine process $P_{B}$ to cause more harm than $P_{Cr}$, when the difference is caused by $P_{A}$, $P_{Cr}$, and $P_{B}$, the correctness of Lemma 4 would be still unaffected.
@@ 5. Summary of Findings
In pursuing the objectives of this work using 3-hop indirect delivery, we found:
1. The number of tolerated link faults / links asynchrony per system size and the actual number of faulty processes.
2. Impossibility of two equal in size largest consensus majority sets operating with bounded delivery.
3. Impossibility of two consensus majority sets of different size with bounded delivery and intersection smaller than a majority.
4. Possibility for $f$-resilience to Byzantine faults implemented with two rounds of exchange of messages.
5. Impossibility of transient nature of links asynchrony and link faults to affect consensus safety, validity, or termination.
6. The same tolerance to max allowed process crashes regardless of whether crashes happened before or during consensus rounds.
7. A way to make leaderless consensus in partial synchrony a non-existing theoretical or practical problem.
8. Possibility for consensus algorithm in partial synchrony with ensured simultaneous validity, safety, and bounded liveness.
@@ 6. Conclusion
We proved the possibility for deterministic Byzantine consensus with synchronous termination in partial synchrony, by showing: i) possibility to tolerate number of asynchronous links, that is a function of the number of faulty processes per system size, with the use of 3-hop epidemic dissemination; and ii) an algorithm that operates with cryptographically authenticated messages and ensures resilience to $f$ Byzantine processes with 2 messaging rounds.
A round with circumvented asynchrony involves exchange of $n^{3}$ messages and lasts 3 times longer than a round in synchrony, where $(f+1)$ rounds are needed for resilience to $f$ Byzantine processes (Dolev and Strong, 1983). Yet a system, resilient to more than 5 Byzantine processes, terminates with the presented solution in partial synchrony faster than with the best known Byzantine algorithm in synchrony. This is paid with 13.56 percent more exchange of data.111 With 100kB initial data per process, 32 bytes hash digest size, and 64 bytes signature size, a system of 11 processes tolerates 5 Byzantine processes, requires 6 communication rounds in synchrony, exchanges 86.12 MB of data with 880 messages, and terminates after 6 times the synchrony delivery bound (Dolev and Strong, 1983). With our algorithm, the same system tolerates the same number of Byzantine processes and terminates within the same time interval in partial synchrony with exchange of 1,980 messages. With the same workload, hash digest size, and signature size, it exchanges 97.80MB of data – just 13.56 percent more.
The presented solution enables the design of non-blocking distributed transactions across decentralised database systems. These transactions are vital for the feasibility of some world-shaking innovations. The International Monetary Fund concept about foreign exchange markets with tokenised currencies and instant settlement (Adrian et al., 2023) showcases their critical role.